# Possible molecules of triple-heavy pentaquarks within the extended local hidden gauge formalism

### v1

@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12335001, 12247101, 11965016 and 11705069, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, the 111 Project under Grant No. B20063, the fundamental Research Funds for the Central Universities under Grant No. lzujbky-2022-sp02, and the project for top-notch innovative talents of Gansu province. This work is also partly supported by the Natural Science Foundation of Changsha under Grants No. kq2208257, the Natural Science Foundation of Hunan province under Grant No. 2023JJ30647, the Natural Science Foundation of Guangxi province under Grant No. 2023JJA110076, and the NSFC under Grant No. 12365019 (C.W.X.).

### v3

@@ Acknowledgements
We thank Professor Eulogio Oset for carefully reading the manuscript and providing valuable comments. This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12335001, 12247101, 11965016 and 11705069, the National Key Research and Development Program of China under Contract No. 2020YFA0406400, the 111 Project under Grant No. B20063, the fundamental Research Funds for the Central Universities under Grant No. lzujbky-2022-sp02, and the project for top-notch innovative talents of Gansu province. This work is also partly supported by the Natural Science Foundation of Changsha under Grants No. kq2208257, the Natural Science Foundation of Hunan province under Grant No. 2023JJ30647, the Natural Science Foundation of Guangxi province under Grant No. 2023JJA110076, and the NSFC under Grant No. 12365019 (C.W.X.).